OpenStudy (anonymous):

Part 1: Solve the inequality −5x + 8 less than or equal to 5x − 2 and show all of your work. You can use <= to represent the less than or equal to symbol. Part 2: Use complete sentences to describe the graph of the solution from Part 1. Be sure to include the type of circle and which direction the shading will be.

OpenStudy (anonymous):

it's different from mine.. @Tazmaniadevil maybe i can work something out hopefully..

OpenStudy (anonymous):

ok hold on

OpenStudy (texaschic101):

-5x + 8 <= 5x - 2 -- subtract 5x from both sides -5x - 5x + 8 <= 5x - 5x - 2 -- combine like terms -10x + 8 <= -2 -- subtract 8 from both sides -10x + 8 - 8 <= -2 - 8 -- combine like terms -10x <= -10 -- divide both sides by -10, and change the inequality sign x >= 1 if there is an equal sign, it is a closed circle no equal sign, open circle (<) less then means shading to the left (>) greater then means shading to the right solution : closed circle on 1, shading to the right any questions ?

OpenStudy (anonymous):

thank you! i understand it much better now @texaschic101

OpenStudy (texaschic101):

glad to help :)